OpenStudy (anonymous):

Solve & PLS HELP!!! PLEASE 1/4t^3-16t=0

OpenStudy (anonymous):

do you mean like... \[\LARGE \frac{1}{4}t^3-16t=0\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

factor out a T first

OpenStudy (anonymous):

then you have the difference of two squares... can you factor it now?

OpenStudy (anonymous):

Not really...

OpenStudy (anonymous):

\[\LARGE t\cdot(\frac14 t^2-16)=0\] now you have t=0 or (1/4)t^2-16=0

OpenStudy (anonymous):

Wah???

OpenStudy (anonymous):

\[\LARGE \frac {t^2}{4}=16 \longrightarrow t^2=64\] \[\LARGE t=\pm \sqrt{64}\] so you have three solutions... t=0 t=8 t=-8

OpenStudy (anonymous):

But isn't it supposed to be two solutions?

OpenStudy (anonymous):

not in this case... only when you have quadratic equation... here you have t^3

OpenStudy (anonymous):

O ok then thanks

OpenStudy (anonymous):

Never mind, Glad to help )